Shawn is planting trees along his driveway, and he has 5 oaks and 5 redwoods to plant in one row. What is the probability that he randomly plants the trees so that all 5 oaks are next to each other and all 5 redwoods are next to each other? Express your answer as a fraction or a decimal number rounded to four decimal places.
Answer #1

Total number of trees 5+5=10

Since 5 trees are same and 5 are same so the possible number of arrangements of 10 trees are

10! 5252

There are two ways in which all 5 oaks are next to each other and all redwoods are next to each other so the required probability is

P = 2 /252 = 0.0079